OpenStudy (anonymous):

Help with simplifying please :) How do you get from 1=sqrt(5)secthetax-2ytantheta to 1= (xsectheta)/2-(ytantheta)/sqrt(5) ??

OpenStudy (anonymous):

not sure is this \[1=\sqrt{5}x\sec(\theta)-2y\tan(\theta)\]

OpenStudy (anonymous):

yep exactly it is in relation to finding the parametric form of a tangent to a hyperbola with P (2sectheta, sqrt5tantheta) if that helps...

OpenStudy (anonymous):

no it doesn't help

OpenStudy (anonymous):

okay well thanks anyway

OpenStudy (anonymous):

looks to me like the left hand side has been devided by \(2\sqrt{5}\) but the right hand side has not

OpenStudy (anonymous):

If anyone else is able to help, i know the final form of the tangent is \[1=(xsec \theta)/2)-((ytan \theta)/\sqrt{5})\] but i am having trouble getting in that form...

OpenStudy (anonymous):

got my left and right backwards

OpenStudy (anonymous):

Yeah- if i could multiply it by a fraction that was equivalent to one it would keep the 1 the same but alter the form of the rhs...just need to figure out that fraction...(not good with fractions)

OpenStudy (anonymous):

hmm the more i look at it, the more it doesn't look the same. one side has been divided, other has not

OpenStudy (anonymous):

I must have a mistake in my working.. back to the drawing board! :) thanks for the help